# Efficient Backdoor Attacks for Deep Neural Networks in Real-world Scenarios

### v1

@@ 6 Limitations and Future Works
Performance degradation in clean-label backdoor attacks. Clean-label backdoor attacks present a significant challenge [69]. As shown in Fig. 7, previous methods exhibit a poor ASR, and our technologies show limited improvement in clean-label backdoor attacks when the poisoning rate is low. In future research, we will investigate the underlying reasons for this situation and explore more efficient attack methods specifically designed for clean-label backdoor attacks.
@@
Application limitations. Our technologies depend on the CLIP model that is pre-trained on natural images, which may limit their applicability to certain domains such as medical images or remote sensing. In such cases, a possible solution is to replace CLIP with a domain-specific pre-trained model, such as MedCLIP [36] for medical images or Satellite [1] for remote sensing, to adapt our methods to the target domain.

### v2

@@ A.16 Limitations and Future Works
In this section, we discuss the limitations of our approach and outline potential future directions for backdoor learning research. Performance Degradation in Clean-label Backdoor Attacks. Clean-label backdoor attacks present a significant challenge Zhao et al. (2020). As shown in Fig. 3, previous methods exhibit a poor ASR, and our technologies show limited improvement in clean-label backdoor attacks when the poisoning rate is low. In future research, we will investigate the underlying reasons for this situation and explore more efficient attack methods specifically designed for clean-label backdoor attacks.
@@
Application Limitations. Our technologies depend on the CLIP model that is pre-trained on natural images, which may limit their applicability to certain domains such as medical images or remote sensing. In such cases, a possible solution is to replace CLIP with a domain-specific pre-trained model, such as MedCLIP Madhawa & Carlomagno (2022) for medical images or Satellite Arto et al. (2021) for remote sensing, to adapt our methods to the target domain.